# LLM2Vec: Large Language Models Are Secretly Powerful Text Encoders

### v1

@@ G.3 Results
6 LLM2Vec (w/o SimCSE) 14.22 $\checkmark$ 4096 55.99 58.42 45.54 87.99 76.63 84.09 29.96 64.80
6 LLM2Vec 14.22 $\checkmark$ 4096 56.05 58.59 45.12 88.18 76.72 83.69 30.66 64.72
@@
Classify a given Amazon customer review text as either counterfactual or not-counterfactual
AmazonPolarityClassification
@@
Classify Amazon reviews into positive or negative sentiment
AmazonReviewsClassification
@@
Classify the given Amazon review into its appropriate rating category
Banking77Classification
@@
Given a online banking query, find the corresponding intents
EmotionClassification
@@
Classify the emotion expressed in the given Twitter message into one of the six emotions: anger, fear, joy, love, sadness, and surprise
ImdbClassification
@@
Classify the sentiment expressed in the given movie review text from the IMDB dataset
MassiveIntentClassification
Given a user utterance as query, find the user intents
MassiveScenarioClassification
Given a user utterance as query, find the user scenarios
MTOPDomainClassification
@@
Classify the intent domain of the given utterance in task-oriented conversation
MTOPIntentClassification
@@
Classify the intent of the given utterance in task-oriented conversation
ToxicConversationsClassif.
Classify the given comments as either toxic or not toxic
TweetSentimentClassification
@@
Classify the sentiment of a given tweet as either positive, negative, or neutral
ArxivClusteringP2P
@@
Identify the main and secondary category of Arxiv papers based on the titles and abstracts
ArxivClusteringS2S
@@
Identify the main and secondary category of Arxiv papers based on the titles
BiorxivClusteringP2P
@@
Identify the main category of Biorxiv papers based on the titles and abstracts
BiorxivClusteringS2S
@@
Identify the main category of Biorxiv papers based on the titles
MedrxivClusteringP2P
@@
Identify the main category of Medrxiv papers based on the titles and abstracts
MedrxivClusteringS2S
@@
Identify the main category of Medrxiv papers based on the titles
RedditClustering
@@
Identify the topic or theme of Reddit posts based on the titles
RedditClusteringP2P
@@
Identify the topic or theme of Reddit posts based on the titles and posts
StackExchangeClustering
@@
Identify the topic or theme of StackExchange posts based on the titles
StackExchangeClusteringP2P
@@
Identify the topic or theme of StackExchange posts based on the given paragraphs
TwentyNewsgroupsClustering
Identify the topic or theme of the given news articles
SprintDuplicateQuestions
Retrieve duplicate questions from Sprint forum
TwitterSemEval2015
@@
Retrieve tweets that are semantically similar to the given tweet
TwitterURLCorpus
@@
Retrieve tweets that are semantically similar to the given tweet
AskUbuntuDupQuestions
Retrieve duplicate questions from AskUbuntu forum
MindSmallReranking
@@
Retrieve relevant news articles based on user browsing history
SciDocsRR
@@
Given a title of a scientific paper, retrieve the titles of other relevant papers
StackOverflowDupQuestions
Retrieve duplicate questions from StackOverflow forum
ArguAna
Given a claim, find documents that refute the claim
ClimateFEVER
@@
Given a claim about climate change, retrieve documents that support or refute the claim
CQADupstackRetrieval
@@
Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
DBPedia
@@
Given a query, retrieve relevant entity descriptions from DBPedia
FEVER
@@
Given a claim, retrieve documents that support or refute the claim
FiQA2018
@@
Given a financial question, retrieve user replies that best answer the question
HotpotQA
@@
Given a multi-hop question, retrieve documents that can help answer the question
MSMARCO
@@
Given a web search query, retrieve relevant passages that answer the query
NFCorpus
@@
Given a question, retrieve relevant documents that best answer the question
NQ
Given a question, retrieve Wikipedia passages that answer the question
QuoraRetrieval
@@
Given a question, retrieve questions that are semantically equivalent to the given question
SCIDOCS
@@
Given a scientific paper title, retrieve paper abstracts that are cited by the given paper
SciFact
@@
Given a scientific claim, retrieve documents that support or refute the claim
Touche2020
@@
Given a question, retrieve detailed and persuasive arguments that answer the question
TRECCOVID
@@
Given a query on COVID-19, retrieve documents that answer the query
STS*
Retrieve semantically similar text.
BUCC/Tatoeba
Retrieve parallel sentences.
SummEval
@@
Given a news summary, retrieve other semantically similar summaries

### v2

@@ G.3 Results
6 Bi + MNTP + Mean 14.22 $\checkmark$ 4096 55.99 58.42 45.54 87.99 76.63 84.09 29.96 64.80
6 Bi + MNTP + SimCSE + Mean 14.22 $\checkmark$ 4096 56.05 58.59 45.12 88.18 76.72 83.69 30.66 64.72
@@
AmazonCounterfactualClassif. Classify a given Amazon customer review text as either counterfactual or not-counterfactual
@@
AmazonPolarityClassification Classify Amazon reviews into positive or negative sentiment
@@
AmazonReviewsClassification Classify the given Amazon review into its appropriate rating category
@@
Banking77Classification Given a online banking query, find the corresponding intents
@@
EmotionClassification Classify the emotion expressed in the given Twitter message into one of the six emotions: anger, fear, joy, love, sadness, and surprise
@@
ImdbClassification Classify the sentiment expressed in the given movie review text from the IMDB dataset
MassiveIntentClassification Given a user utterance as query, find the user intents
MassiveScenarioClassification Given a user utterance as query, find the user scenarios
@@
MTOPDomainClassification Classify the intent domain of the given utterance in task-oriented conversation
@@
MTOPIntentClassification Classify the intent of the given utterance in task-oriented conversation
ToxicConversationsClassif. Classify the given comments as either toxic or not toxic
@@
TweetSentimentClassification Classify the sentiment of a given tweet as either positive, negative, or neutral
@@
ArxivClusteringP2P Identify the main and secondary category of Arxiv papers based on the titles and abstracts
@@
ArxivClusteringS2S Identify the main and secondary category of Arxiv papers based on the titles
@@
BiorxivClusteringP2P Identify the main category of Biorxiv papers based on the titles and abstracts
@@
BiorxivClusteringS2S Identify the main category of Biorxiv papers based on the titles
@@
MedrxivClusteringP2P Identify the main category of Medrxiv papers based on the titles and abstracts
@@
MedrxivClusteringS2S Identify the main category of Medrxiv papers based on the titles
@@
RedditClustering Identify the topic or theme of Reddit posts based on the titles
@@
RedditClusteringP2P Identify the topic or theme of Reddit posts based on the titles and posts
@@
StackExchangeClustering Identify the topic or theme of StackExchange posts based on the titles
@@
StackExchangeClusteringP2P Identify the topic or theme of StackExchange posts based on the given paragraphs
TwentyNewsgroupsClustering Identify the topic or theme of the given news articles
SprintDuplicateQuestions Retrieve duplicate questions from Sprint forum
@@
TwitterSemEval2015 Retrieve tweets that are semantically similar to the given tweet
@@
TwitterURLCorpus Retrieve tweets that are semantically similar to the given tweet
AskUbuntuDupQuestions Retrieve duplicate questions from AskUbuntu forum
@@
MindSmallReranking Retrieve relevant news articles based on user browsing history
@@
SciDocsRR Given a title of a scientific paper, retrieve the titles of other relevant papers
StackOverflowDupQuestions Retrieve duplicate questions from StackOverflow forum
ArguAna Given a claim, find documents that refute the claim
@@
ClimateFEVER Given a claim about climate change, retrieve documents that support or refute the claim
@@
CQADupstackRetrieval Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
@@
DBPedia Given a query, retrieve relevant entity descriptions from DBPedia
@@
FEVER Given a claim, retrieve documents that support or refute the claim
@@
FiQA2018 Given a financial question, retrieve user replies that best answer the question
@@
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question
@@
MSMARCO Given a web search query, retrieve relevant passages that answer the query
@@
NFCorpus Given a question, retrieve relevant documents that best answer the question
NQ Given a question, retrieve Wikipedia passages that answer the question
@@
QuoraRetrieval Given a question, retrieve questions that are semantically equivalent to the given question
@@
SCIDOCS Given a scientific paper title, retrieve paper abstracts that are cited by the given paper
@@
SciFact Given a scientific claim, retrieve documents that support or refute the claim
@@
Touche2020 Given a question, retrieve detailed and persuasive arguments that answer the question
@@
TRECCOVID Given a query on COVID-19, retrieve documents that answer the query
STS* Retrieve semantically similar text.
BUCC/Tatoeba Retrieve parallel sentences.
@@
SummEval Given a news summary, retrieve other semantically similar summaries